OpenStudy (anonymous):

find the midpoint of the line segment connecting the points (-5,3) and (6,-9)

OpenStudy (anonymous):

See to find the midpoints simply add the x coordinates and then divide it by 2.. And do the same for y coordinates..

OpenStudy (anonymous):

\[Midpoint(x, y) = (\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2})\]

OpenStudy (anonymous):

to find the midpoint of two coordinates \[(x _{1},y _{1}) and (x _{2},y _{2})\] the midpoint is given by \[M(\frac{ x _{1},+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 })\]

OpenStudy (anonymous):

so 1/2,-3?

OpenStudy (anonymous):

Yep you are right.. Good..

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

You are welcome dear..